OpenStudy (anonymous):

Which has the same value as (4 + 3i)2?

OpenStudy (tommo_lcfc):

4 + 3i squared or 2(4+3i)?

OpenStudy (tommo_lcfc):

(4+3i)(4+3i) = 16 + 24i + 9i^2 = 16 + 24i + 9(-1) = 24i + 7